OpenStudy (anonymous):

Five more than a number tripled is half the number. What is the number? A. 6 B. –2 C. –1 D. 1

OpenStudy (whpalmer4):

a number: x a number tripled: 3x 5 more than a number tripled: 3x + 5 half the number: x/2 can you solve it?

OpenStudy (anonymous):



OpenStudy (anonymous):

idk how to do it

OpenStudy (anonymous):

just read the the word and express it in mathematical notations

OpenStudy (texaschic101):

5 more then a number triples is half the number 3x + 5 = 1/2x 5 = 1/2x - 3x 5 = 1/2x - 6/2x 5 = - 5/2x 5/(-5/2) =x 5 * -2/5 = x - 10/5 = x -2 = x check.. 3x + 5 = 1/2x 3(-2) + 5 = 1/2(-2) -6 + 5 = -2/2 -1 = -1 (correct) ANSWER: -2

OpenStudy (whpalmer4):

I parsed the words for you: \[3x+5 = \frac{x}{2}\]To solve this, I would multiply everything by 2 \[6x + 10 = x\]Subtract x from both sides \[5x + 10 = 0\]subtract 10 from both sides \[5x = -10\]Divide both sides by 5 \[x = -2\]